# Controllability score for linear time-invariant systems on an infinite time horizon

### v1

@@ III-B Infinite-horizon scaled controllability Gramian
Let $C_{1}\in\mathbb{C}^{\ell_{1}\times\ell_{1}},C_{2}\in\mathbb{C}^{\ell_{2}\times\ell_{2}}$, and $P\in\mathbb{C}^{\ell_{1}\times\ell_{2}}$. If $\mu_{1}+\mu_{2}\neq 0$ holds for any eigenvalues $\mu_{1}$ of $C_{1}$ and $\mu_{2}$ of $C_{2}$, then
$C_{1}Y+YC_{2}+P=0$
has a unique solution $Y\in\mathbb{C}^{\ell_{1}\times\ell_{2}}$.
@@
where we set $\alpha=m_{-}+m_{0}^{(1)}+\cdots+m_{0}^{(\ell-1)},\ \beta=m_{-}+m_{0}^{(1)}+\cdots+m_{0}^{(\ell)}$ to simplify the indexing.
@@ III-C Reformulation of controllability scoring problems
Since the uniqueness of the optimal solutions is equivalent as well, we obtain the same result as in [23, Theorem 5] and [22, Theorem 1]. In Appendix E, we summarize existing results on the uniqueness of the optimal solutions and present new results.
@@ IV Extension to infinite-horizon settings
In what follows, we define
$W_{\infty,i}\coloneq W_{\infty}(e_{i}e_{i}^{\top}),$
for which
@@ IV-A Problem formulation
We refer to the optimal solution to Problem 5 as the VCS on an infinite time horizon, or simply the infinite-horizon VCS. Similarly, we refer to the optimal solution to Problem 6 as the AECS on an infinite time horizon, or simply the infinite-horizon AECS. In the case where $A$ is Hurwitz, Problems 5 and 6 coincide with the controllability scoring problems proposed in [23].
Since $W(p;T)\to W_{\infty}(p)$, Problem 5 is formally obtained from Problem 3 by letting $T\to\infty$, and it depends on the whole matrix $W_{\infty}(p)$. On the other hand, since $D(T)^{-1}\to\begin{pmatrix}I_{n_{-}}\\ &O\end{pmatrix}$, formally letting $T\to\infty$ in Problem 4 yields the objective function of Problem 6, which depends only on $W_{-}(p)$; consequently, the constraint is posed only on $W_{-}(p)$.
@@
Problem 5 imposes a constraint that ensures controllability of the entire system, whereas Problem 6 requires controllability only of the stable modes. Consequently, while the infinite-horizon VCS enforces controllability of the full system, the infinite-horizon AECS may render the entire system uncontrollable.
@@ IV-C Uniqueness analysis
We discuss the uniqueness of the optimal solutions to Problems 5 and 6. The uniqueness of the optimal solution is important for two reasons.
@@
Indeed, by taking $Q=\dfrac{1}{\sqrt{2}}\begin{pmatrix}1&-1\\ 1&1\end{pmatrix}$, we obtain $Q^{-1}AQ=\operatorname{diag}(1,-1)$, whose eigenvalues are $1$ and 1-1. Therefore, by Theorem 7, Problems 1 and 2 each admit a unique optimal solution for all $T>0$.
@@
We partition $Q^{-1}e_{i}$ into blocks in following two ways:
$Q^{-1}e_{i}=\begin{pmatrix}q_{1,i}\\ \vdots\\ q_{m,i}\end{pmatrix}=\begin{pmatrix}q_{-,i}\\ q_{0,i}\\ q_{+,i}\end{pmatrix}$.
Furthermore, let $q^{\prime}_{k,i}$ denote the $n_{k}$-th (i.e., the last) entry of $q_{k,i}$. For $\ell=1,\ldots,m_{0}^{\prime}$, we define
$q_{0,i}^{(\ell)}\coloneq(q^{\prime}_{\alpha+1,i},\ldots,q^{\prime}_{\beta,i})^{\top}\in\mathbb{C}^{m_{0}^{(\ell)}},$
where we set $\alpha=m_{-}+m_{0}^{(1)}+\cdots+m_{0}^{(\ell-1)},\ \beta=m_{-}+m_{0}^{(1)}+\cdots+m_{0}^{(\ell)}$ to simplify the indexing. Then, the following theorems hold.
@@ V Algorithms
In this section, we describe an algorithm for solving Problems 5 and 6. In formulating Problems 5 and 6, we employed the Jordan canonical form; however, a direct use of the Jordan canonical form may lead to numerical instability [8]. As shown below, under the following assumption, the optimal solutions to Problems 5 and 6 can be computed in a numerically stable manner via the real Schur decomposition and the solutions of Sylvester equations.
@@
and the diagonal entries satisfy
$\ell_{ii}=-\sum_{j\neq i}\ell_{ij}.$
When considering graph Laplacians, we restrict attention to positive edge weights whenever an edge exists. For undirected graphs, the Laplacian matrix $L$ is symmetric, whereas for directed graphs, $L$ is generally not symmetric.
Let $L$ be the Laplacian matrix of a directed graph. Then, A=LA=-L satisfies Assumption 2. In fact, from the definition, the matrix $A=(a_{ij})$ satisfies
$a_{ij}\geq 0\ (i\neq j),\quad a_{ii}=-\sum_{j\neq i}a_{ij}\leq 0.$
It follows from Geršhgorin’s theorem [27, Theorem 1.1] that all eigenvalues of $A$ have nonpositive real parts, and that zero is the only eigenvalue that may lie on the imaginary axis. Furthermore, $\operatorname{rank}(A)=\operatorname{rank}(A^{2})$ holds [6, Proposition 12]. This implies that the zero eigenvalue of $A$ is semisimple. $\Box$
First, we perform a block diagonalization using the real Schur decomposition and the solution of a Sylvester equation. Since this approach does not require the Jordan canonical form, it is numerically stable. By the real Schur decomposition, we can find an orthogonal matrix $R^{(1)}\in\mathbb{R}^{n\times n}$ such that
@@
Since no two of the matrices $A_{-},A_{0}$, and $A_{+}$ share a common eigenvalue, 14, 15, and 16 can be solved successively and admit a unique solution from Lemma 1. Define
$R^{(2)}\coloneq\begin{pmatrix}I_{n_{-}}\\ R^{(2)}_{0,-}&I_{n_{0}}\\ R^{(2)}_{+,-}&R^{(2)}_{+,0}&I_{n_{+}}\end{pmatrix}$
and set $R=R^{(1)}R^{(2)}$. Then, by using
@@
$A_{-}V_{-,i}+V_{-,i}A_{-}^{\top}+q_{-,i}r_{-,i}^{\top}=0,$ (20)
$A_{+}V_{+,i}+V_{+,i}A_{+}^{\top}-q_{+,i}r_{+,i}^{\top}=0.$ (21)
Moreover, let $V_{0,i}\coloneq r_{0,i}r_{0,i}^{\top}$. Finally, for $p=(p_{1},\ldots,p_{n})^{\top}\in\mathbb{R}^{n}$, we define
@@
Since $A$ is block-diagonalized as in 18, the Jordan canonical forms of $A_{-},A_{0}$, and $A_{+}$ are given by $J_{,}J_{0}$, and $J_{+}$, respectively. Thus, there exist invertible matrices $S_{-}\in\mathbb{C}^{n_{-}\times n_{-}},\ S_{0}\in\mathbb{C}^{n_{0}\times n_{0}}$, and $S_{+}\in\mathbb{C}^{n_{+}\times n_{+}}$ such that $S_{-}^{-1}A_{-}S_{-}=J_{-},\ S_{0}^{-1}A_{0}S_{0}=J_{0}$, and $S_{+}^{-1}A_{+}S_{+}=J_{+}$. Moreover, letting
$S\coloneq\begin{pmatrix}S_{-}\\ &S_{0}\\ &&S_{+}\end{pmatrix}$,
we obtain $S^{-1}(R^{-1}AR)S=J$; hence $S_{-},S_{0}$ and $S_{+}$ can be chosen so that $RS=Q$.
As defined in Section IV-C,
$\begin{pmatrix}q_{-,i}\\ q_{0,i}\\ q_{+,i}\end{pmatrix}=Q^{-1}e_{i}=S^{-1}\begin{pmatrix}r_{-,i}\\ r_{0,i}\\ r_{+,i}\end{pmatrix}$.
Then, $r_{-,i}=S_{-}q_{-,i},\ r_{0,i}=S_{0}q_{0,i}$, and $r_{+,i}=S_{+}q_{+,i}$ hold.
@@
$J_{-}W_{-,i}+W_{-,i}J_{-}^{*}+r_{-,i}r_{-,i}^{*}=0,$ (22)
$(-J_{+})W_{+,i}+W_{+,i}(-J_{+}^{*})+r_{+,i}r_{+,i}^{*}=0,$ (23)
respectively. Multiplying 22 from the left by $S_{-}$ and from the right by $S_{-}^{*}$, and using the uniqueness of solutions to 20, we obtain $V_{-,i}=S_{-}W_{-,i}S_{-}^{*}$. Similary, it follows from 21 and 23 that $V_{+,i}=S_{+}W_{+,i}S_{+}^{*}$. Since $W_{0,i}$ reduces to $q_{0,i}q_{0,i}^{*}$ under Assumption 2, we obtain $V_{0,i}=S_{0}W_{0,i}S_{0}^{*}$. Therefore,
@@
$=\operatorname{tr}\left(R^{(1)}\begin{pmatrix}V_{-}(p)^{-1}\\ &O\end{pmatrix}(R^{(1)})^{\top}\right)$
$=\operatorname{tr}\left(V_{-}(p)^{-1}\right),$
@@
Once $V_{\mathrm{l},i},V_{\mathrm{i},i}$, and $V_{\mathrm{r},i}$ have been computed, Problems 7 and 8 take the same form as the Problems 1 and 2. Hence, we can employ Algorithms 2 and 3, which are the same as those in [23]. Here, $f$ is the objective function, and $\Pi_{\Delta}$ is a Euclidean projection onto $\Delta$.
@@
$g(p)\coloneq-\log\det V(p),$
$h(p)\coloneq\operatorname{tr}\left(V_{-}(p)^{-1}\right).$
@@ VII-A Summary
In this paper, we introduced a scaled controllability Gramian and used it to reformulate conventional controllability scoring problems into forms that can be solved in a numerically stable manner even when the system matrix $A$ is non-Hurwitz. By further taking the limit $T\to\infty$ in these reformulated problems, we formulated infinite-horizon controllability scoring problems. Under suitable assumptions, we proved that the finite-horizon controllability scores converge to the infinite-horizon ones and that the infinite-horizon controllability scores are unique. This proves a principled choice of the terminal time, namely $T=\infty$.
Although the formulations of infinite-horizon controllability scoring problems rely on the Jordan canonical form of $A$, we also showed that, under appropriate assumptions, their optimal solutions can be computed using the Schur decomposition and the solution of Sylvester equations, and we presented a corresponding numerical algorithm. In the numerical experiments, we computed infinite-horizon controllability scores for the same Laplacian dynamics as in [22] and observed that finite-horizon controllability scores indeed converge to infinite-horizon ones as $T\to\infty$.
@@ Appendix A Proof of Theorem 1
Let $C_{1}\in\mathbb{C}^{\ell_{1}\times\ell_{1}},C_{2}\in\mathbb{C}^{\ell_{2}\times\ell_{2}}$, and $R\in\mathbb{C}^{\ell_{1}\times\ell_{2}}$. If $\operatorname{Re}(\mu_{1}+\mu_{2})<0$ holds for any eigenvalues $\mu_{1}$ of $C_{1}$ and $\mu_{2}$ of $C_{2}$, then
$\int_{0}^{\infty}\mathrm{e}^{C_{1}t}P\mathrm{e}^{C_{2}t}\mathrm{d}t$
converges, and it is the unique solution $Y\in\mathbb{C}^{\ell_{1}\times\ell_{2}}$ satisfying
$C_{1}Y+YC_{2}+P=0.$
@@
From 4, 5, 26, and 27, the $(i,j)$-block of $W(B;T)$ can be written as
@@
In the following, we analyze the block representation 28 by distinguishing the following nine cases:
@@
Noting that $D_{i}(T)=I_{n_{i}}$ and $D_{j}(T)=I_{n_{j}}$, 28 reduces to
$\int_{0}^{T}\mathrm{e}^{J_{i}t}\widetilde{B}_{i,j}\mathrm{e}^{J_{j}^{*}t}\mathrm{d}t.$
Since $\operatorname{Re}(\lambda_{i}),\operatorname{Re}(\lambda_{j})<0$, it follows from Lemma 4 that the $(i,j)$-block converges, as $T\to\infty$, to the unique solution $W_{-}^{(i,j)}(B)\in\mathbb{C}^{n_{i}\times n_{j}}$ satisfying
$J_{i}W_{-}^{(i,j)}(B)+W_{-}^{(i,j)}(B)J_{j}^{*}+\widetilde{B}_{i,j}=0.$
@@
Noting that $D_{i}(T)=I_{n_{i}}$, 28 reduces to
@@
Noting that $D_{i}(T)=I_{n_{i}}$ and $D_{j}(T)=\mathrm{e}^{J_{j}T}$, 28 reduces to
@@
There exists a constant $c\geq 0$ such that
$\left\|{\mathrm{e}^{J_{i}t}}\right\|\leq ct^{n_{i}-1}\mathrm{e}^{\operatorname{Re}(\lambda_{i})t},$
$\left\|{\mathrm{e}^{J_{j}^{*}t}}\right\|\leq ct^{n_{j}-1}\mathrm{e}^{\operatorname{Re}(\lambda_{j})t},$
$\left\|{\widetilde{B}_{i,j}}\right\|\leq c$
for all $t\geq 0$. Consequently, we obtain
@@
Since
$\mathrm{e}^{J_{i}t}=\sum_{k=0}^{n_{i}-1}\mathrm{e}^{\lambda_{i}t}\dfrac{t^{k}}{k!}N_{n_{i}}^{k},$
we obtain
@@
for some constant $c\geq 0$. Consequently, the summand of 29 converges, as $T\to\infty$, to
@@
Noting that $D_{j}(T)=\mathrm{e}^{J_{j}T}$, we can rewrite 28 as
@@
Noting that $D_{i}(T)=\mathrm{e}^{J_{i}T}$ and $D_{j}(T)=\mathrm{e}^{J_{j}T}$, 28 reduces to
$\int_{0}^{T}\mathrm{e}^{J_{i}(t-T)}\widetilde{B}_{i,j}\mathrm{e}^{J_{j}^{*}(t-T)}\mathrm{d}t$
$=\int_{0}^{T}\mathrm{e}^{-J_{i}t}\widetilde{B}_{i,j}\mathrm{e}^{-J_{j}^{*}t}\mathrm{d}t.$
Since $\operatorname{Re}(\lambda_{i}),\operatorname{Re}(\lambda_{j})>0$, it follows from Lemma 4 that the integral of the right-hand side converges, as $T\to\infty$, to the unique solution $W_{+}^{(i,j)}(B)\in\mathbb{C}^{n_{i}\times n_{j}}$ satisfying

### v2

@@ III-B Infinite-horizon scaled controllability Gramian
Let $C_{1}\in\mathbb{C}^{\ell_{1}\times\ell_{1}},C_{2}\in\mathbb{C}^{\ell_{2}\times\ell_{2}}$, and $P\in\mathbb{C}^{\ell_{1}\times\ell_{2}}$. If $\mu_{1}+\mu_{2}\neq 0$ holds for any eigenvalues $\mu_{1}$ of $C_{1}$ and $\mu_{2}$ of $C_{2}$, then $C_{1}Y+YC_{2}+P=O$ has a unique solution $Y\in\mathbb{C}^{\ell_{1}\times\ell_{2}}$.
@@
where we set $\alpha=m_{-}+m_{0}^{(1)}+\cdots+m_{0}^{(\ell-1)},\ \beta=\alpha+m_{0}^{(\ell)}$ to simplify the indexing.
@@ III-C Reformulation of controllability scoring problems
Since the uniqueness of the optimal solutions is equivalent as well, we obtain the same result as in [24, Theorem 5] and [23, Theorem 1]. In Appendix E, we summarize existing results on the uniqueness of the optimal solutions and present new results.
@@ IV Extension to infinite-horizon settings
In what follows, we define $W_{\infty,i}\coloneq W_{\infty}(e_{i}e_{i}^{\top})$, for which
@@ IV-A Problem formulation
We refer to the optimal solution to Problem 5 as the VCS on an infinite time horizon, or simply the infinite-horizon VCS. Since $W(p;T)\to W_{\infty}(p)$ as $T\to\infty$, Problem 5 is obtained from Problem 3 by taking the limit $T\to\infty$, and its objective depends on the full matrix $W_{\infty}(p)$.
@@
We refer to the optimal solution to Problem 6 as the AECS on an infinite time horizon, or simply the infinite-horizon AECS. As $T\to\infty$, we have $D(T)^{-1}\to\begin{pmatrix}I_{n_{-}}\\ &O\end{pmatrix}$, and Theorem 1 then shows that the objective in Problem 4 converges to that of Problem 6. In particular, the latter depends only on $W_{-}(p)$, and consequently the constraint is imposed only on $W_{-}(p)$.
If $A$ is Hurwitz, then Problems 5 and 6 coincide with the controllability scoring problems proposed in [24].
@@
Problem 5 imposes a constraint that ensures controllability of the entire system, whereas Problem 6 requires controllability only of the stable modes. Consequently, while the infinite-horizon VCS enforces controllability of the full system, the infinite-horizon AECS may render the entire system uncontrollable. We illustrate this phenomenon with a numerical example in Section VI.
@@ IV-C Uniqueness analysis
We discuss uniqueness of the optimal solutions to Problems 5 and 6, which is important for two reasons.
@@
Indeed, by taking $Q=\dfrac{1}{\sqrt{2}}\begin{pmatrix}1&-1\\ 1&1\end{pmatrix}$, we obtain $Q^{-1}AQ=\operatorname{diag}(1,-1)$, whose eigenvalues are $1$ and 1-1. Therefore, $\Theta^{\prime}=\emptyset$ in Theorem 7 of Appendix E, and hence, by Theorem 7, Problems 1 and 2 each admit a unique optimal solution for all $T>0$.
@@
We partition $Q^{-1}e_{i}$ into blocks in following two ways: $Q^{-1}e_{i}=\begin{pmatrix}q_{1,i}\\ \vdots\\ q_{m,i}\end{pmatrix}=\begin{pmatrix}q_{-,i}\\ q_{0,i}\\ q_{+,i}\end{pmatrix}$. Furthermore, let $q^{\prime}_{k,i}$ denote the $n_{k}$-th (i.e., the last) entry of $q_{k,i}$. For $\ell=1,\ldots,m_{0}^{\prime}$, we define $q_{0,i}^{(\ell)}\coloneq(q^{\prime}_{\alpha+1,i},\ldots,q^{\prime}_{\beta,i})^{\top}\in\mathbb{C}^{m_{0}^{(\ell)}}$, where we set $\alpha=m_{-}+m_{0}^{(1)}+\cdots+m_{0}^{(\ell-1)},\ \beta=\alpha+m_{0}^{(\ell)}$ to simplify the indexing. Then, the following theorems hold.
@@ V Algorithms
As discussed in Section I, the observations in [23] suggest choosing $T$ sufficiently large in Problems 1 and 2, in which case the resulting problems can be solved using existing algorithms [24]. On the other hand, Theorems 2 and 3 show that Problems 5 and 6 yield the limiting values of these finite-horizon controllability scores. Moreover, beyond yielding the limit, Problems 5 and 6 admit a numerically stable computation—even for unstable systems—via the scaled-Gramian reformulation, whereas the direct finite-horizon computation can suffer from numerical instability when $T$ is large. Finally, solving Problems 5 and 6 is computationally more efficient than solving Problems 1 and 2 for large values of $T$, as described below.
In formulating Problems 5 and 6, we employed the Jordan canonical form; however, a direct use of the Jordan canonical form may lead to numerical instability [8]. As shown below, under the following assumption, the optimal solutions to Problems 5 and 6 can be computed in a numerically stable manner via the real Schur decomposition and solutions of Sylvester equations.
@@
and the diagonal entries satisfy $\ell_{ii}=-\sum_{j\neq i}\ell_{ij}$. When considering graph Laplacians, we restrict attention to positive edge weights whenever an edge exists. For undirected graphs, the Laplacian matrix $L$ is symmetric, whereas for directed graphs, $L$ is generally not symmetric.
Let $L$ be the Laplacian matrix of a directed graph. Then, A=LA=-L satisfies Assumption 2. In fact, from the definition, the matrix $A=(a_{ij})$ satisfies $a_{ij}\geq 0\ (i\neq j)$ and $a_{ii}=-\sum_{j\neq i}a_{ij}\leq 0$. It follows from Geršhgorin’s theorem [28, Theorem 1.1] that all eigenvalues of $A$ have nonpositive real parts, and that zero is the only eigenvalue that may lie on the imaginary axis. Furthermore, $\operatorname{rank}(A)=\operatorname{rank}(A^{2})$ holds [6, Proposition 12]. This implies that the zero eigenvalue of $A$ is semisimple. $\Box$
First, we perform a block diagonalization using the real Schur decomposition and the solution of Sylvester equations. Since this approach does not require the Jordan canonical form, it is numerically stable. By the real Schur decomposition, we can find an orthogonal matrix $R^{(1)}\in\mathbb{R}^{n\times n}$ such that
@@
Since no two of the matrices $A_{-},A_{0}$, and $A_{+}$ share a common eigenvalue, 14, 15, and 16 can be solved successively and admit a unique solution from Lemma 1. Define $R^{(2)}\coloneq\begin{pmatrix}I_{n_{-}}\\ R^{(2)}_{0,-}&I_{n_{0}}\\ R^{(2)}_{+,-}&R^{(2)}_{+,0}&I_{n_{+}}\end{pmatrix}$ and set $R=R^{(1)}R^{(2)}$. Then, by using
@@
$A_{-}V_{-,i}+V_{-,i}A_{-}^{\top}+q_{-,i}r_{-,i}^{\top}=O,$ (20)
$(-A_{+})V_{+,i}+V_{+,i}(-A_{+}^{\top})+q_{+,i}r_{+,i}^{\top}=O.$ (21)
Moreover, let
$V_{0,i}\coloneq r_{0,i}r_{0,i}^{\top}.$ (22)
Finally, for $p=(p_{1},\ldots,p_{n})^{\top}\in\mathbb{R}^{n}$, we define
@@
Since $A$ is block-diagonalized as in 18, the Jordan canonical forms of $A_{-},A_{0}$, and $A_{+}$ are given by $J_{,}J_{0}$, and $J_{+}$, respectively. Thus, there exist invertible matrices $S_{-}\in\mathbb{C}^{n_{-}\times n_{-}},\ S_{0}\in\mathbb{C}^{n_{0}\times n_{0}}$, and $S_{+}\in\mathbb{C}^{n_{+}\times n_{+}}$ such that $S_{-}^{-1}A_{-}S_{-}=J_{-},\ S_{0}^{-1}A_{0}S_{0}=J_{0}$, and $S_{+}^{-1}A_{+}S_{+}=J_{+}$. Moreover, letting $S\coloneq\begin{pmatrix}S_{-}\\ &S_{0}\\ &&S_{+}\end{pmatrix}$, we obtain $S^{-1}(R^{-1}AR)S=J$; hence $S_{-},S_{0}$ and $S_{+}$ can be chosen so that $RS=Q$.
As defined in Section IV-C, $\begin{pmatrix}q_{-,i}\\ q_{0,i}\\ q_{+,i}\end{pmatrix}=Q^{-1}e_{i}=S^{-1}\begin{pmatrix}r_{-,i}\\ r_{0,i}\\ r_{+,i}\end{pmatrix}$. Then, $r_{-,i}=S_{-}q_{-,i},\ r_{0,i}=S_{0}q_{0,i}$, and $r_{+,i}=S_{+}q_{+,i}$ hold.
@@
$J_{-}W_{-,i}+W_{-,i}J_{-}^{*}+r_{-,i}r_{-,i}^{*}=O,$ (23)
$(-J_{+})W_{+,i}+W_{+,i}(-J_{+}^{*})+r_{+,i}r_{+,i}^{*}=O,$ (24)
respectively. Multiplying 23 from the left by $S_{-}$ and from the right by $S_{-}^{*}$, and using the uniqueness of solutions to 20, we obtain $V_{-,i}=S_{-}W_{-,i}S_{-}^{*}$. Similary, it follows from 21 and 24 that $V_{+,i}=S_{+}W_{+,i}S_{+}^{*}$. Since $W_{0,i}$ reduces to $q_{0,i}q_{0,i}^{*}$ under Assumption 2, we obtain $V_{0,i}=S_{0}W_{0,i}S_{0}^{*}$. Therefore,
@@
$=\operatorname{tr}\left(R^{(1)}\begin{pmatrix}V_{-}(p)^{-1}\\ &O\end{pmatrix}(R^{(1)})^{\top}\right)=\operatorname{tr}\left(V_{-}(p)^{-1}\right),$
@@
Once $V_{-,i},V_{0,i}$, and $V_{+,i}$ have been computed, Problems 7 and 8 take the same form as the Problems 1 and 2. Hence, we can employ Algorithms 2 and 3, which are the same as those in [24]. Here, $f$ is the objective function, and $\Pi_{\Delta}$ is a Euclidean projection onto $\Delta$.
@@
$g(p)\coloneq-\log\det V(p),\quad h(p)\coloneq\operatorname{tr}\left(V_{-}(p)^{-1}\right).$
@@
When $T$ is large but finite, computing $\widetilde{W}(e_{i}e_{i}^{\top};T)$ or $W(e_{i}e_{i}^{\top};T)$ via numerical integration leads to a high computational cost that increases with $T$. In contrast, matrices $V_{i}$ that play the same role as the infinite-horizon scaled controllability Gramians $W_{\infty,i}$ can be computed by solving the Lyapunov equations 20 and 21 together with computing 22, whose computational costs are independent of $T$. Moreover, Lyapunov equations can be solved efficiently in an approximate manner using algorithms such as CF-ADI [15], making the infinite-horizon formulation advantageous from a computational perspective.
@@ VII-A Summary
In this paper, we introduced a scaled controllability Gramian and used it to reformulate conventional controllability scoring problems into forms that can be solved in a numerically stable manner even when the system matrix $A$ is non-Hurwitz. By further taking the limit $T\to\infty$ in these reformulated problems, we formulated infinite-horizon controllability scoring problems. Under suitable assumptions, we proved that the finite-horizon controllability scores converge to the infinite-horizon ones and that the infinite-horizon controllability scores are unique. This provides a principled choice of the terminal time, namely $T=\infty$.
Although the formulations of infinite-horizon controllability scoring problems rely on the Jordan canonical form of $A$, we also showed that, under appropriate assumptions, their optimal solutions can be computed using the Schur decomposition and the solution of Sylvester equations, and we presented a corresponding numerical algorithm. Moreover, the infinite-horizon formulation is more advantageous than the finite-horizon one from a computational perspective. In the numerical experiments, we computed infinite-horizon controllability scores for the same Laplacian dynamics as in [23] and observed that finite-horizon controllability scores indeed converge to infinite-horizon ones as $T\to\infty$.
@@ Appendix A Proof of Theorem 1
Let $C_{1}\in\mathbb{C}^{\ell_{1}\times\ell_{1}},C_{2}\in\mathbb{C}^{\ell_{2}\times\ell_{2}}$, and $P\in\mathbb{C}^{\ell_{1}\times\ell_{2}}$. If $\operatorname{Re}(\mu_{1}+\mu_{2})<0$ holds for any eigenvalues $\mu_{1}$ of $C_{1}$ and $\mu_{2}$ of $C_{2}$, then $\int_{0}^{\infty}\mathrm{e}^{C_{1}t}P\mathrm{e}^{C_{2}t}\mathrm{d}t$ converges, and it is the unique solution $Y\in\mathbb{C}^{\ell_{1}\times\ell_{2}}$ satisfying $C_{1}Y+YC_{2}+P=0$.
@@
From 4, 5, 27, and 28, the $(i,j)$-block of $W(B;T)$ can be written as
@@
In the following, we analyze the block representation 29 by distinguishing the following nine cases:
@@
Noting that $D_{i}(T)=I_{n_{i}}$ and $D_{j}(T)=I_{n_{j}}$, 29 reduces to $\int_{0}^{T}\mathrm{e}^{J_{i}t}\widetilde{B}_{i,j}\mathrm{e}^{J_{j}^{*}t}\mathrm{d}t$. Since $\operatorname{Re}(\lambda_{i}),\operatorname{Re}(\lambda_{j})<0$, it follows from Lemma 4 that the $(i,j)$-block converges, as $T\to\infty$, to the unique solution $W_{-}^{(i,j)}(B)\in\mathbb{C}^{n_{i}\times n_{j}}$ satisfying
$J_{i}W_{-}^{(i,j)}(B)+W_{-}^{(i,j)}(B)J_{j}^{*}+\widetilde{B}_{i,j}=O.$
@@
Noting that $D_{i}(T)=I_{n_{i}}$, 29 reduces to
@@
Noting that $D_{i}(T)=I_{n_{i}}$ and $D_{j}(T)=\mathrm{e}^{J_{j}T}$, 29 reduces to
@@
There exists a constant $c\geq 0$ such that $\left\|{\mathrm{e}^{J_{i}t}}\right\|\leq ct^{n_{i}-1}\mathrm{e}^{\operatorname{Re}(\lambda_{i})t}$, $\left\|{\mathrm{e}^{J_{j}^{*}t}}\right\|\leq ct^{n_{j}-1}\mathrm{e}^{\operatorname{Re}(\lambda_{j})t}$, and $\left\|{\widetilde{B}_{i,j}}\right\|\leq c$ for all $t\geq 0$. Consequently, we obtain
@@
Since $\mathrm{e}^{J_{i}t}=\sum_{k=0}^{n_{i}-1}\mathrm{e}^{\lambda_{i}t}\dfrac{t^{k}}{k!}N_{n_{i}}^{k}$, we obtain
@@
for some constant $c\geq 0$. Consequently, the summand of 30 converges, as $T\to\infty$, to
@@
Noting that $D_{j}(T)=\mathrm{e}^{J_{j}T}$, we can rewrite 29 as
@@
Noting that $D_{i}(T)=\mathrm{e}^{J_{i}T}$ and $D_{j}(T)=\mathrm{e}^{J_{j}T}$, 29 reduces to $\int_{0}^{T}\mathrm{e}^{J_{i}(t-T)}\widetilde{B}_{i,j}\mathrm{e}^{J_{j}^{*}(t-T)}\mathrm{d}t=\int_{0}^{T}\mathrm{e}^{-J_{i}t}\widetilde{B}_{i,j}\mathrm{e}^{-J_{j}^{*}t}\mathrm{d}t$. Since $\operatorname{Re}(\lambda_{i}),\operatorname{Re}(\lambda_{j})>0$, it follows from Lemma 4 that the integral of the right-hand side converges, as $T\to\infty$, to the unique solution $W_{+}^{(i,j)}(B)\in\mathbb{C}^{n_{i}\times n_{j}}$ satisfying